# On the exact solutions of a two-dimensional hydrogen atom in a constant magnetic field

### v1

@@ Abstract
We analyse the exact solutions to some quantum-mechanical models derived recently by means of a generalization of the Frobenius method. In particular, we carry out an interpretation of the polynomial solutions obtained for the two-dimensional hydrogen atom in a constant magnetic field by comparing the exact solutions with numerical results provided by the Rayleigh-Ritz method.
@@ 1 Introduction
Recently, Bildstein and Grabowski[13] proposed a generalization of the Frobenius method and applied it to quantum-mechanical models with polynomial potentials in one, two and three dimensions. They obtained analytical solutions for most of them. In particular, they considered in detail the two-dimensional hydrogen atom in a constant magnetic field already studied by Le et al[12] and somewhat earlier by Taut[14]. The purpose of this paper is the interpretation of the exact QS solutions for this model.

### v2

@@ Abstract
We discuss the exact polynomial solutions for the two-dimensional hydrogen atom in a constant magnetic field already studied earlier by other authors. In order to provide a suitable meaning for such solutions we compare them with numerical results provided by the Rayleigh-Ritz method.
@@ 1 Introduction
Recently, Bildstein and Grabowski[13] proposed a generalization of the Frobenius method and applied it to quantum-mechanical models with polynomial potentials in one, two and three dimensions. They obtained analytical solutions for most of them. In particular, they considered in detail the two-dimensional hydrogen atom in a constant magnetic field already studied by Le et al[12] and somewhat earlier by Taut[14]. In particular, Turbiner and Escobar Ruiz discussed the existence of a hidden algebra. The purpose of this paper is the interpretation of the exact QS solutions for this model.